OpenStudy (anonymous):

Choose the factored form of the given equation. 6x^2 + 2x - 8 = 0 (6x- 8)(x+ 1) = 0 (6x+ 1)(x- 8) = 0 (6x+ 8)(x- 1) = 0

OpenStudy (anonymous):

ok all the first terms work right? each has \(6x\times x=6x^2\) and all the last terms work because each one gies \(-8\) so the only question is which one gives \(2x\)

OpenStudy (anonymous):

= 6x+ 8)(x- 1) = 0 be cause x1=-8 and x2=1 (Viet)

OpenStudy (anonymous):

first one \[6x-8x=-2x\] no second one \[-48x+1=-47x\] no third one \[-6x+8x=2x\] yes